Shift report 20000812 08:00-16:00

From: http account (http@pii3.brahms.bnl.gov)
Date: Sat Aug 12 2000 - 15:58:34 EDT

  • Next message: Konstantin Olchanski: "Avoid using pii4.brahms.bnl.gov"

    ------------------------------------------------------------
    Date: Sat Aug 12 15:58:33 2000, Shift: 08:00-16:00
    Supervisor: RD, with: ZIS
    Log book#4 pages 62 - 74
    
    Shift Summary:
    ------------------------------------------------------------
    Machine status: Great stores! One from previous shift and a second
    that started at 11:00 and is still going. At the beginning of the store
    we had 8x10^9 in blue and 6x10^9 in yellow. The best rate 
    for BRAHMS was some 60Hz in scaler 1 (ZDC trigger)
    
    2. Achievements and runs:
        run 2202, 2203, 2205  D1 on D5 and D2 off
        run 2206, 2207  D5 at 4Kgauss
        run 2209, 2216  MRS moved at 60 degrees D2 on in 2216
        run 2217, 2218  D5 on
        run 2219, 2220, 2221 T1 gating grid studies
        run 2222 continue MRS program
    
    3 Detector prob.
        T2 readout got bad; trigger receiver on RDB is suspected
    culprit, may need swapping of boards
    
         We couldn't mode MRS to 45 degrees because we saw a 
    yellow caution cable tying BB stand to TPM1 stand. We will
    have an access at end of this store and cut it.
    
         T1 gating grid studies are confusing: with the gate closed
    we see more tracks compared to runs where the gating is pulsing.
    
         At the beginning of the store (60 Hz interactions) T1
    anodes tripped every ~10 min. Now the rate is as low as 1hr.
    T2A also tripped 3 times during the shift.
    
    Flemming has something to add:
    
    The trigger setup was continued. The width (time) of the
    trigger one Left and Right pulses were increaed by a few nsec to about 7 
    nsec. This should ensure that all valid trigger 6 (two large tubes in L,R)
    always sets trigger 1, that is less central.
    
    The timing of H1 ,but not H2 gate, was changed. This was a leftover effort
    from the trigger time alignement last weekend. The timing
    looks ok on scope (a group appears at about 70 nsec). This should be checked 
    by FS TOF people. by looking at data (following 2222).
    
    The Rhic blue clock was re-aligned with the ZDC Y.B (it show up
    now as scaler 9 in pet page)
    
    The sync scaler input (1 sec clock had been dis-connected is back
    in scaler display.)
    
    DB opening ZDC ADC gate for high gain (120 nsec) and moved gate
    another 10 nsec earlier. For low gain gate was moved 20 nsec (page 67)
     
    Several signals added to trigger scalers that are read out event per event.
      
    ------------------------------------------------------------
    



    This archive was generated by hypermail 2b29 : Sat Aug 12 2000 - 16:02:00 EDT